solving systems Using Elimination : 2x-3y=-11
3x+2y=29

Answer by mananth(16946)

2x-3y= -11 .............1
3x+2y=29 .............2
Eliminate y
multiply (1)by 2
Multiply (2) by 3
4x-6y=-22
9x+6y= 87
Add the two equations
13x=65
/13
x= 5
plug value of x in (1)
2x-3y=-11
10-3y=-11
-3y=-11-10
-3y=-21
y=7